OpenStudy (anonymous):

Which of the following is equal to \(\sqrt(18x^7~y^7\)? 2x^5y^4 9x^6y^5 3x^5y^4sqrt(2) 3x^3y^4sqrt(2x)

OpenStudy (anonymous):

3x^3y^3 squareroot 2xy is what i got

OpenStudy (anonymous):

so 3x^3y^4 sqrt 2x so D

OpenStudy (anonymous):

how'd u get 3x^3y^3 squareroot 2xy

OpenStudy (loser66):

still need help?

OpenStudy (anonymous):

sorry for the late reply i was eating dinner, yeah I really have no clue on how to do this

OpenStudy (loser66):

work with me is hard, ready?

OpenStudy (anonymous):

i'm ready

OpenStudy (loser66):

under the square root, you have 3 terms, number, x and y. We solve one by one now number 18 = 9*2 ok? and 9 = \(3^2\) so, 18 = \(3^2 *2\) got this?

OpenStudy (loser66):

now, you tell me 2\(\div\)2 =? remainder =?

OpenStudy (anonymous):

1?

OpenStudy (loser66):

remainder?

OpenStudy (anonymous):

oh 0 there is no remainder

OpenStudy (loser66):

ok, let it there, we will be back later, I give you a new concept. surely you will master this kind of problem 

OpenStudy (loser66):

give me the answer of 5\(\div\)2= remainder

OpenStudy (anonymous):

a^2.5 5/2 - 2.5 ?

OpenStudy (loser66):

be patient, not that,, just answer me, I spend time to tutor you step by step , if you jump, you cannot get the method.

OpenStudy (anonymous):

ok but 5/2 does equal 2.5 right?

OpenStudy (loser66):

yes, but no, because I ask you for quotient and remainder

OpenStudy (anonymous):

i'm confused what do you mean by quotient? I thought the remainder is what you were left with which is 2.5 so what is a quotient (sorry for being so stupid :c )

OpenStudy (loser66):